How do you find the other endpoint of the line segment with the given endpoint and midpoint
Endpoint:(8,9), midpoint:(-10,4)

Answer by jim_thompson5910(35256)

(8+x)/2 = -10

8+x = 2*(-10)

8+x = -20

x = -20 - 8

x = -28

-------------------------------------------------------

(9+y)/2 = 4

9+y = 2*4

9+y = 8

y = 8 - 9

y = -1

-------------------------------------------------------

So the other endpoint is (-28, -1)